# scenarioparampadding=8pt, margin-top=3pt, margin-bottom=3pt, border-style=none, background-color=quotebackground

### v1

@@ Abstract.
In the mental health domain, Large Language Models (LLMs) offer promising new opportunities, though their inherent complexity and low controllability have raised questions about their suitability in clinical settings. We present MindfulDiary, a mobile journaling app incorporating an LLM to help psychiatric patients document daily experiences through conversation. Designed in collaboration with mental health professionals (MHPs), MindfulDiary takes a state-based approach to safely comply with the experts’ guidelines while carrying on free-form conversations. Through a four-week field study involving 28 patients with major depressive disorder and five psychiatrists, we found that MindfulDiary supported patients in consistently enriching their daily records and helped psychiatrists better empathize with their patients through an understanding of their thoughts and daily contexts. Drawing on these findings, we discuss the implications of leveraging LLMs in the mental health domain, bridging the technical feasibility and their integration into clinical settings.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06submissionid: 3044ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Natural language interfaces
@@ 1. Introduction
Journals serve as a written record of an individual’s past events, thoughts, and feelings, allowing genuine expression  (Travers, 2011; Ullrich and Lutgendorf, 2002). Prior work has shown the advantages of journaling in clinical mental health contexts, as journals frequently capture patients’ daily experiences, symptoms, and other contextual data that are challenging to gather during brief hospital visits (Figueiredo et al., 2020; Zhu et al., 2016). Furthermore, these patient journals can enhance mental health professionals’ (MHPs) comprehension of their patients’ conditions, leading to improved treatment quality (Wu et al., 2020). However, writing about one’s past feelings and thoughts can be a complex process because people differ in their ability to understand, identify, and verbalize their emotions (Salovey and Mayer, 1990). In addition, patient under psychotherapy struggle with constructing a narrative and understanding their past (Donnelly and Murray, 1991; Pennebaker and Seagal, 1999).
@@
Researchers in the field of Human-Computer Interaction (HCI) have shown that chatbots can help individuals articulate and share their daily experiences. For instance, chatbots to elicit people’s self-disclosure can ease the process of emotional expression by providing a safe and supportive environment for individuals to share their experiences and emotions (Park et al., 2021; Lee et al., 2020b, a; Cameron et al., 2019). Furthermore, a machine’s inherent trait of not showing fatigue can make people more confident to share their stories truthfully and comfortably (Kim et al., 2018; Park et al., 2021).
The recent advancements in large language models (LLMs) have added to the understanding and generation of natural conversations. Unlike traditional chatbot (Abd-alrazaq et al., 2019) were unable to support conversation outside pre-defined context (Wang et al., 2021; Dsouza et al., 2019), chatbot driven by LLMs can offer dynamic responses on diverse topics (Bae et al., 2022; Roller et al., 2021; Brown et al., 2020). However, care needs to be taken when leveraging LLMs, particularly in mental health contexts, as the inherent challenge of LLMs in uncertainty in control could result in generating unintended or inaccurate responses (Farhat, 2023; Wang et al., 2021; Korngiebel and Mooney, 2021).
@@
In this work, we present MindfulDiary, a mobile conversational AI to support patients in journaling their daily experiences and thoughts through conversations. MindfulDiary incorporates LLMs to generate a response, prompting patients differently according to the conversational phase. The conversation records are automatically summarized and presented on a clinician dashboard so MHPs can obtain insights about the patient. As a multi-disciplinary research team, which included HCI researchers, AI engineers, and psychiatrists, we iteratively designed MindfulDiary and conducted a four-week field study with 28 psychiatric patients diagnosed with major depressive disorder (MDD). Through this study, we found that the versatility, narrative-building capability, and diverse perspectives provided by MindfulDiary assisted patients in consistently enriching their daily records. Furthermore, MindfulDiary supported patients in overcoming the challenges of detailed record-keeping and expression, often hindered by feelings of apathy and cognitive burdens. Notably, the enriched records from MindfulDiary provided MHPs with deeper insights, enhancing their understanding and empathy toward their patients. Based on the findings, we discuss the potential of LLMs’ conversational abilities to aid in patient record-keeping within clinical mental health settings and suggestions for their responsible integration into these environments.
Our work contribute the following:
@@
    We present MindfulDiary, an LLM-driven journal designed to document psychiatric patients’ daily experiences through naturalistic conversations, designed in collaboration with MHPs.
@@
    From a four-week field study involving 28 patients and five psychiatrists, we provide empirical findings on how MindfulDiary supported patients in keeping their daily logs and assisted psychiatrists in monitoring and comprehending patient states.
@@
    Drawing from the design process of MindfulDiary and the findings of our deployment study, we outline key considerations and share lessons learned for the design and implementation of LLM’s conversational capability in clinical mental health settings.
@@ 2. Related Work
In this section, we cover related work in two parts: (1) journal as a patient-generated health data, and (2) conversational agent for mental health.
2.1. Journal as a Patient-generated Health Data in Clinical Setting
@@
PGHD—defined as “health-related data, such as health history, symptoms, biometric readings, treatment history, lifestyle choices, and other pertinent details, created, recorded, or gathered by patients” (Shapiro et al., 2012)—has increasingly become an essential tool in clinical settings to capture authentic, real-time insights into patients‘ health. Studies have shown that PGHD can enhance communication between patients and MHPs and offer contextual information about patients, thereby heightening MHPs’ awareness of patient health outside regular clinical visits (Lordon et al., 2020; Cohen et al., 2016; Nundy et al., 2014).
@@
Within the mental health domain, PGHD range from structured mental health assessments (e.g., anxiety, depression) to more unstructured data, including mood-related symptoms and social interactions tracking (e.g., social media use, number of calls)(Austin et al., 2020; Demiris et al., 2019; Nghiem et al., 2023). Our study places particular emphasis on journals as a PGHD method in clinical mental health settings. Journaling refers to “the practice of writing down one‘s symptoms and other information related to one‘s daily life to discuss them during clinical appointments” (Figueiredo et al., 2020; Zhu et al., 2016). Using a journal as a PGHD can particularly be useful in mental health contexts as it can offer rich, self-documented insights, which could improve MHPs‘ understanding of their patients (Wu et al., 2020).
However, while the benefits of journaling are clear, the act itself is not always straightforward. Many people struggle with starting their entries, sticking to consistent journaling routines, and structuring their reflections (Travers, 2011). Further, writing about emotions and past experiences can be intricate, as individuals vary in their capacity to recognize, interpret, and articulate their feelings (Salovey and Mayer, 1990; Lane et al., 1990; Luminet et al., 2004). For some, especially in psychotherapy, crafting a narrative that describes one’s life journey can be a challenging process (Donnelly and Murray, 1991; Pennebaker and Seagal, 1999).
While previous research has extensively established the significance of employing journals as PGHD within clinical contexts, it is essential to acknowledge the inherent complexities tied to the journaling process. This becomes especially critical when contemplating its application within the specific demographic undergoing psychotherapy. With this background, our research explores the potential of LLMs in assisting the journaling process within the clinical mental health setting.
2.2. Conversational Agent for Mental Health
The field of artificial intelligence (AI) has proposed some significant innovations in medical settings, such as aiding clinical decision-making and diagnosis (Magrabi et al., 2019; Beam and Kohane, 2016). However, AIs have rarely been integrated into mental health contexts, partly because the field has relied on nuanced techniques and subjective data, such as patient relationships and emotion observation (Graham et al., 2019; Gabbard and Crisp-Han, 2017). Recently, there has been a growing technological presence of Natural Language Processing (NLP) applications for supporting mental health, as NLP’s capability for understanding human language holds a significant promise in this domain (Calvo et al., 2017; Demner-Fushman et al., 2009).
@@ 2.2. Conversational Agent for Mental Health
Conversational agents, also known as chatbots, have particularly stood out in the mental health domain (Abd-Alrazaq et al., 2021; Abd-alrazaq et al., 2019). Studies have demonstrated the potential of chatbots in facilitating different types of therapy, such as cognitive behavioral therapy (Fitzpatrick et al., 2017; Inkster et al., 2018; Fulmer et al., 2018), expressive writing (Park et al., 2021), behavioral reinforcement (Inkster et al., 2018), and solution-focused therapy (Fulmer et al., 2018). Prior work has also shown that chatbots could ease the burden of disclosing sensitive information. Studies indicated that individuals may feel more comfortable communicating with chatbots because of the social stigma involved in communicating with human beings (Lee et al., 2020b; Park et al., 2021; Lee et al., 2020a). Furthermore, these approaches can help overcome temporal and spatial constraints, offering mental health support that is accessible at any time and anywhere (Cameron et al., 2019).
Recently, chatbots powered by LLMs have gained attention (Wang et al., 2021). Trained on vast amounts of text data, they demonstrate exceptional performance and versatility in natural open-domain conversation, surpassing previous NLP models in understanding intricate language patterns (Radford et al., 2019; Tanana et al., 2021). Yet, many chatbots used for mental health therapy still employ a retrieval-based approach (Abd-alrazaq et al., 2019; Abd-Alrazaq et al., 2019). The basic principle of retrieval-based chatbots is to search for an appropriate response from a database based on input from users. Since the responses and expected flow that the system can provide are already hard-coded in the database, chatbots exhibit consistent and predictable behavior (Dsouza et al., 2019). However, if a user gives input that does not exist in the database, the chatbot may fail to respond. Thus, these systems cannot flexibly respond to unforeseen situations. In contrast, a generative approach driven by LLM has made it possible to develop chatbots that can produce human-like and contextualized outputs for even unforeseen input, enabling open-domain dialogues (Bae et al., 2022).
@@
However, LLMs, employing a neural network known as a “transformer” to generate outputs, come with inherent challenges tied to their architecture (Vaswani et al., 2017). One key issue is the explainability of the model output: it’s challenging to discern how this ‘black box’ model interprets a given input prompt. As a result, designers struggle to predict both the LLM’s understanding of the input and the subsequent messages it might produce (Liu et al., 2023). For instance, a chatbot leveraging GPT-2 for mental health therapy occasionally generated non-word sentences or produced more negative outputs than positive ones (Wang et al., 2021). Replika, an LLM-based application intended for mental well-being, has occasionally displayed harmful content and exhibited inconsistent conversational styles, undermining its role as a long-term companion (Ma et al., 2023).
Considering the inherent challenges of LLM and the sensitivity of the mental health domain, we must reflect carefully on how and to what extent the conversational capabilities of these models should be utilized. In this study, we emphasize the need to consider the opportunities and limitations of LLMs, as well as important aspects of the mental health domain, including safety, sensitivity, and experts’ empirical knowledge. With this in mind, we designed MindfulDiary’s state-based chatbot using both rule-based and generation-based approaches to safe compliance with experts’ guidelines while facilitating free-form conversations.

### v2

# MindfulDiary: Harnessing Large Language Model to Support Psychiatric Patients’ Journaling

@@ Abstract.
Large Language Models (LLMs) offer promising opportunities in mental health domains, although their inherent complexity and low controllability elicit concern regarding their applicability in clinical settings. We present MindfulDiary, an LLM-driven journaling app that helps psychiatric patients document daily experiences through conversation. Designed in collaboration with mental health professionals, MindfulDiary takes a state-based approach to safely comply with the experts’ guidelines while carrying on free-form conversations. Through a four-week field study involving 28 patients with major depressive disorder and five psychiatrists, we examined how MindfulDiary facilitates patients’ journaling practice and clinical care. The study revealed that MindfulDiary supported patients in consistently enriching their daily records and helped clinicians better empathize with their patients through an understanding of their thoughts and daily contexts. Drawing on these findings, we discuss the implications of leveraging LLMs in the mental health domain, bridging the technical feasibility and their integration into clinical settings.
@@
journalyear: 2024copyright: rightsretainedconference: Proceedings of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613904.3642937isbn: 979-8-4007-0330-0/24/05submissionid: 3044ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Natural language interfaces
Image of MindfulDiary’s interface split into two sections. On the left, the patient interface displays a journal session screen which aids user reflection through prompt question generation and journal-style summarization. On the right, the expert interface showcases a desktop dashboard layout, providing professionals with insights into journal interactions, key events and emotions, as well as the trend of PHQ-9 scores.
@@ 1. Introduction
Journals serve as a written record of an individual’s past events, thoughts, and feelings, allowing genuine expression (Travers, 2011; Ullrich and Lutgendorf, 2002). Journaling helps people describe experiences and express emotions related to both negative (Pennebaker, 1997; Pennebaker and Chung, 2011) and positive experiences (e.g., growth potential) (Baikie et al., 2012; Guastella and Dadds, 2009), thereby reducing stress, anxiety, and depression. Prior work has shown the advantages of journaling in clinical mental health contexts, as journals frequently capture patients’ daily experiences, symptoms, and other contextual data that are challenging to gather during brief hospital visits (Figueiredo et al., 2020; Zhu et al., 2016). Furthermore, these patient journals can enhance mental health professionals (MHPs) comprehension of their patient’s conditions, leading to improved treatment quality (Wu et al., 2020). However, writing about one’s past feelings and thoughts can be a complex process because people differ in their ability to understand, identify, and verbalize their emotions (Salovey and Mayer, 1990). In addition, patient under psychotherapy struggle with constructing a narrative and understanding their past (Donnelly and Murray, 1991; Pennebaker and Seagal, 1999).
@@
Conversational AIs, or chatbots, have the potential as an alternative form of journaling, easing the collection of personal data. Researchers in the field of Human-Computer Interaction (HCI) have shown that chatbots can help individuals articulate and share their daily experiences. For instance, chatbots to elicit people’s self-disclosure can ease the process of emotional expression by providing a safe and supportive environment for individuals to share their experiences and emotions (Park et al., 2021; Lee et al., 2020b, a; Cameron et al., 2019). Furthermore, a machine’s inherent trait of not showing fatigue can make people more confident to share their stories truthfully and comfortably (Kim et al., 2018; Park et al., 2021). However, existing chatbot prototypes have commonly employed rule-based or retrieval-driven approaches (Abd-alrazaq et al., 2019), which have limited capability of generating versatile responses following up serendipitous topics during conversation (Huang et al., 2020; Lee et al., 2019; Jo et al., 2023). This trend presents missed opportunities and a lack of understanding regarding conversational AIs that assist with journaling by suggesting, questioning, and empathizing based on the user’s diverse experiences.
The recent achievement of Natural Language Processing in large language models (LLMs) opened up new opportunities for bootstrapping chatbots that can carry on more naturalistic conversation (Bae et al., 2022; Jo et al., 2023; Wei et al., 2024; Roller et al., 2021; Brown et al., 2020). Their capabilities accelerated the development of chatbots in varied topics that can benefit from open-ended conversation, such as regular check-up calls (Jo et al., 2023; Bae et al., 2022), personal health tracking (Wei et al., 2024), and personal events and emotions (Seo et al., 2023). Despite such opportunities, LLMs’ inherent uncertainty in control of response generation calls for precautions to handle unintended or inaccurate responses (Farhat, 2023; Wang et al., 2021; Korngiebel and Mooney, 2021; Jo et al., 2023). If applied to clinical and mental health domains, LLM’s behaviors should be designed in collaboration with domain experts regarding the relevance and safety of responses.
In this work, we present a case of collaborative design, development, and evaluation of an LLM-infused conversational AI system designed to facilitate the self-reflection of patients and communication with MHPs. We designed and developed MindfulDiary (Figure 1), which consists of (1) a mobile conversational AI with which patients can converse about daily experiences and thoughts and (2) a web dashboard that allows MHPs to review their patients’ dialogue history with the AI. MindfulDiary incorporates LLMs to generate a response, prompting patients differently according to the conversational phase. The conversation records are automatically summarized and presented on a clinician dashboard so MHPs can obtain insights about the patient.
@@
As a multi-disciplinary research team, which included HCI researchers, AI engineers, and psychiatrists, we iteratively designed MindfulDiary and conducted a four-week field study involving 28 psychiatric patients diagnosed with major depressive disorder (MDD) and five psychiatrists who care for them. During the study, the patients freely used MindfulDiary to record daily conversations, and the psychiatrists used the clinician dashboard during regular clinical visits. Through this study, we found that the versatility, narrative-building capability, and diverse perspectives provided by MindfulDiary assisted patients in consistently enriching their daily records. Furthermore, MindfulDiary supported patients in overcoming the challenges of detailed record-keeping and expression, often hindered by feelings of apathy and cognitive burdens. The psychiatrists reported that enhanced records provided by MindfulDiary offered a more nuanced understanding of their patients, fostering empathy. In addition, MindfulDiary supplemented their consultation by eliciting candid thoughts from patients that may be invasive to be asked by the MHPs.
The key contributions of this work are:
@@
    Design and development of MindfulDiary, an LLM-driven journal designed to document psychiatric patients’ daily experiences through naturalistic conversations, designed in collaboration with MHPs.
@@
    Empirical findings from a four-week field study involving 28 patients and five psychiatrists, demonstrating how MindfulDiary supported patients in keeping their daily logs and assisted psychiatrists in monitoring and comprehending patient states. We also explore how MindfulDiary enhances the quality of patient-provider communication, emphasizing the role of LLMs in prompting deeper self-exploration, which can be instrumental in clinical settings.
@@
    Implications for designing and instrumenting LLM-infused conversational AIs in clinical mental health settings.
@@ 2. Related Work
In this section, we cover related work in three parts: (1) Journaling for self-reflection and mental Health, (2) Journaling as patient-generated health data in clinical settings, and (3) Conversational agents for mental health.
2.1. Journaling for Self-Reflection and Mental Health
Journaling—recording various personal matters ranging from observations and travels to overall daily experiences and thoughts (Riddell et al., 2020)—takes place in various forms, including letter-style entries reminiscent of Anne Frank’s diary or more traditional prose. By narrating personal experiences and innermost feelings in a journal, people discover insights upon reflecting on their past events (Ullrich and Lutgendorf, 2002; Travers, 2011), thereby improving mental wellness and quality of life (Pennebaker, 1997; Pennebaker and Chung, 2011; Baikie et al., 2012; Utley and Garza, 2011). Researchers have found that journaling impacts mental health by fostering emotional inhibition (Pennebaker, 1985), cognitive processing (Harber et al., 1992), and “freeing up” cognitive load (Klein and Boals, 2001; Ramirez and Beilock, 2011).
Beyond traditional pen and paper, research in HCI has explored how technology can augment a journaling method in articulating past emotions and experiences (e.g.(Elsden et al., 2016; Kalnikaite and Whittaker, 2012; Hodges et al., 2006, 2006; Kocielnik et al., 2018; Bakker and Rickard, 2018; Apple Inc., 2024)). Early studies in journaling focused on how technology can assist users in better documenting past events and experiences from life-logging perspectives (Elsden et al., 2016; Kalnikaite and Whittaker, 2012; Hodges et al., 2006). For example, SenseCam has been proposed as a wearable ubiquitous computing device that utilizes camera sensor data to not only record but also reflect on the wearer’s daily life (Hodges et al., 2006). Subsequently, various attempts have been made to utilize contextual data and cues such as activity levels (Kocielnik et al., 2018), mood (Bakker and Rickard, 2018), location, and photos (Apple Inc., 2024) in journaling context.
Furthermore, HCI researchers have investigated ways to improve a journaling method to encourage users to express themselves in a manner that is more comfortable (Gonzales et al., 2010), engaging (Ma et al., 2017), and honest (Park et al., 2021). For example, Park et al. found that conversational agents create a social-like environment, encouraging self-reflection and enhancing expressive writing (Park et al., 2021). Furthermore, social support from online communities has been found to increase user participation in journaling activities (Ma et al., 2017). Gonzales et al. (Gonzales et al., 2010) proposed an approach to mitigate the discomfort of revisiting negative memories using the sound generated from the data for more pleasurable (Gonzales et al., 2010).
Our work extends this line of research to enhance the journaling experience with technology. Specifically, we explore the potential of interactive dialogue with LLMs as a new format for journaling, allowing psychiatric patients to explore and reflect on their past experiences and emotions. To this end, we leverage a conversational agent as a complementary tool for assisting patients by providing prompt questions to engage users in deeper and more detailed documentation and reflection (Kocielnik et al., 2018). We particularly investigate how our approach plays a role in clinical settings as patient-generated health data, which we will cover in the following.
2.2. Journaling as Patient-Generated Health Data in Clinical Settings
@@
Patient-Generated Health Data (PGHD)—defined as “health-related data, such as health history, symptoms, biometric readings, treatment history, lifestyle choices, and other pertinent details, created, recorded, or gathered by patients” (Shapiro et al., 2012)—has increasingly become an essential tool in clinical settings to capture authentic, real-time insights into patients‘ health. Studies have shown that PGHD can enhance communication between patients and MHPs and offer contextual information about patients, thereby heightening MHPs’ awareness of patient health outside regular clinical visits (Lordon et al., 2020; Cohen et al., 2016; Nundy et al., 2014). For instance, photo journaling improved patient-provider communication for the management and treatment of irritable bowel syndrome by facilitating more effective discussions during treatment (Chung, 2017).
@@
Within the mental health domain, PGHD range from structured mental health assessments (e.g., anxiety, depression) to more unstructured data, including mood-related symptoms and social interactions tracking (e.g., social media use, number of calls)(Austin et al., 2020; Demiris et al., 2019; Nghiem et al., 2023). Patient diaries or journals–“instructing the patient to write down one’s symptoms and other information related to one’s daily life to discuss them during clinical appointments” (Figueiredo et al., 2020; Zhu et al., 2016)–can particularly be useful in mental health contexts as it can offer rich, self-documented insights, which could improve MHPs‘ understanding of their patients (Wu et al., 2020). Despite the benefits of patient journaling, people often struggle with starting their entries, sticking to consistent journaling routines, and structuring their reflections (Travers, 2011). Further, writing about emotions and past experiences can be intricate, as individuals vary in their capacity to recognize, interpret, and articulate their feelings (Salovey and Mayer, 1990; Lane et al., 1990; Luminet et al., 2004). For some, especially in psychotherapy, crafting a narrative that describes one’s life journey can be a challenging process (Donnelly and Murray, 1991; Pennebaker and Seagal, 1999).
In this work, we aim to lower the barrier of journaling for patients by allowing them to carry on casual conversations with an AI instead of plain open-ended text writing. We demonstrate the potential of these dialogues as a source of PGHD in clinical settings to facilitate individual self-reflection and enhance communication between patients and MHPs.
2.3. Conversational Agents for Mental Health
@@
The field of AI has proposed significant innovations in medical settings, such as aiding clinical decision-making and diagnosis (Magrabi et al., 2019; Beam and Kohane, 2016). In a mental health domain, Natural Language Processing (NLP) techniques have been widely applied to venues that demand human language comprehension and generation for patient support and treatment (Calvo et al., 2017; Demner-Fushman et al., 2009). Since caring mental health often involves counseling, conversational agents, also known as chatbots, have particularly stood out in the mental health domain (Abd-Alrazaq et al., 2021; Abd-alrazaq et al., 2019). Studies have demonstrated the potential of chatbots in facilitating different types of therapy, such as cognitive behavioral therapy (Fitzpatrick et al., 2017; Inkster et al., 2018; Fulmer et al., 2018), expressive writing (Park et al., 2021), behavioral reinforcement (Inkster et al., 2018), and solution-focused therapy (Fulmer et al., 2018). Prior work has also shown that chatbots could ease the burden of disclosing sensitive information. Studies indicated that individuals may feel more comfortable communicating with chatbots because of the social stigma involved in communicating with human beings (Lee et al., 2020b; Park et al., 2021; Lee et al., 2020a). Furthermore, these approaches can help overcome temporal and spatial constraints, offering mental health support that is accessible at any time and anywhere (Cameron et al., 2019).
Early mental health therapy chatbots predominantly relied upon rule-based or retrieval-based approaches (Abd-alrazaq et al., 2019; Abd-Alrazaq et al., 2019). While these approaches provide a high level of control over the conversational flow, they fall short of carrying on open-ended conversations; that is, the chatbots neither respond to serendipitous topics nor offer versatile responses that are outside the scope of design (Huang et al., 2020; Gao et al., 2018; Lee et al., 2019). In the mental health context, these limitations may impact the quality of the chatbot’s caring behaviors as their messages tend to be general (Park et al., 2021; Huang et al., 2020). Recent LLMs have suggested a new paradigm of bootstrapping chatbots  (Wang et al., 2021; Wei et al., 2024). LLM-driven chatbots tend to produce human-like, context-aware responses for unseen topics (Wei et al., 2024; OpenAI, Inc, 2022). Consequently, LLM-driven chatbots excel in facilitating open-domain dialogues and engaging in unscripted interactions (Bae et al., 2022), offering flexibility and adaptability in conversations (Radford et al., 2019; Tanana et al., 2021), especially in complex scenarios like mental health support. For instance, GPT-3.5 demonstrates empathetic traits, such as recognizing emotions and offering emotionally supportive replies in various situations, predominantly in healthcare settings (Sorin et al., 2023). In certain cases, these models appear to have potential in tasks that involve empathy, showing promising results when compared to humans (Ayers et al., 2023; Elyoseph et al., 2023). Given these early findings, LLM-driven chatbots in the public health sector may offer some support in alleviating emotional burden and loneliness among isolated individuals, though this area is still under exploration (Jo et al., 2023). Additionally, the integration of LLMs with human efforts in creating mental health peer support messages could potentially lead to more empathetic conversations (Sharma et al., 2023).
@@
However, LLMs suffer from inherent challenges tied to their transformer-based architecture (Vaswani et al., 2017). One key issue is the explainability of the model output: it is challenging to discern how this ‘black box’ model interprets a given input prompt. As a result, designers struggle to predict both the LLM’s understanding of the input and the subsequent messages it might produce (Liu et al., 2023). For instance, a chatbot leveraging GPT-2 for mental health therapy occasionally generated non-word sentences or produced more negative outputs than positive ones (Wang et al., 2021). Replika, an LLM-driven application intended for mental well-being, has occasionally displayed harmful content and exhibited inconsistent conversational styles, undermining its role as a long-term companion (Ma et al., 2023).
These findings highlight two implications for LLM-infused systems in clinical settings. First, human oversight and moderation are critical when using LLMs in clinical settings (Wang et al., 2021). To ensure safe and ethical instrumentation of LLMs for mental health patients, this work involves MHPs to reflect domain experts’ perspectives (Torous et al., 2023) in developing an LLM-driven chatbot system. Second, it is necessary to enhance the controllability of an LLM (Jo et al., 2023; Zamfirescu-Pereira et al., 2023) to better follow the intended conversational design. As one exemplar approach to enhance LLM’s controllability, yet not a conversational domain, AI Chains break down a complex task into multiple, simpler sub-tasks so that individual LLM inferences can work in better reliability (Wu et al., 2022). Similarly, in our system, the chatbot operates using state-based prompting, where a model prompt contains the instruction focused solely on the current state, which is part of an overarching conversational protocol. By simplifying the model prompts, we intended the generated responses to safely and reliably comply with the MHP’s guidelines for interacting with the patients.
In summary, our work leverages LLMs in two key components: First, we use an LLM to power a chatbot for patient journaling. Second, the clinician dashboard incorporates LLMs for various NLP tasks, such as text summarization and classification, to visualize summarized insights (e.g.(Arakawa et al., 2023; Gebreegziabher et al., 2023)) that are noteworthy for MHP’s treatment. Through this collaborative approach involving clinicians, our work explores a unique design space for designing deploying LLM-driven chatbots in the mental health domain, aiming to improve communication between patients and providers by facilitating the recording of daily experiences, which act as a bridge for a better understanding of patients.